one of the games at a carnival involves trying to ring a bell with a ball by hitting a lever that propels the ball into the air. The height of the ball is modeled by the equation h(t)= -16t^2+39t. if the bell is 25ft above the ground, will it be hit by the ball?
Answer by nerdybill(7384)

one of the games at a carnival involves trying to ring a bell with a ball by hitting a lever that propels the ball into the air. The height of the ball is modeled by the equation h(t)= -16t^2+39t. if the bell is 25ft above the ground, will it be hit by the ball?
.
h(t)= -16t^2+39t
The "vertex" will be the max height of the ball.
t (at vertex) = -b/(2a)
t = -39/(2*(-16))
t = -39/(-32)
t = 1.21875 seconds
.
Max height at that time is:
h(1.21875)= -16(1.21875)^2+39(1.21875)
h(1.21875)= -16(1.4853515625)+39(1.21875)
h(1.21875)= -23.765625+47.53125
h(1.21875)= 23.765625
.
Answer: NO